OpenStudy (anonymous):

Anita can exercise for 20 minutes on her lunch break and 30 minutes after work on her 5 work days. She exercises 35 minutes on the weekend. How many hours per week does Anita exercise?

OpenStudy (ivettef365):

what is 20 + 30

OpenStudy (anonymous):

50

OpenStudy (harsimran_hs4):

exercise on 1 working day = 20 + 30 = 50 mins exercise of 5 working days = 5 * (50) = 250 mins total = 250 + 35 = 285 min 60 min = 1h can you do now ?

OpenStudy (ivettef365):

ok 50 minutes a day times 5 days

OpenStudy (anonymous):

250

OpenStudy (ivettef365):

ok 250 + 35 minutes of the weekend

OpenStudy (anonymous):

285

OpenStudy (ivettef365):

ok so the total minutes exercise is 285, now to convert in hours there are 60 minutes in an hour so divide 285/60

OpenStudy (anonymous):

4.75...

OpenStudy (ivettef365):

ok, so total hours are 4 with 45 minutes